OpenStudy (anonymous):

Problem: Steve likes to entertain friends at parties with “wire tricks.” Suppose he takes a piece of wire 60 inches long and cuts it into two pieces. Steve takes the first piece of wire and bends it into the shape of a perfect circle. He then proceeds to bend the second piece of wire into the shape of a perfect square. Where should Steve cut the wire so that the total area of the circle and square combined is as small as possible? What is this minimal area? What should Steve do if he wants the combined area to be as large as possible? NEED HELP!!

OpenStudy (anonymous):

Area of a circle = pi*r^2 Circumference of a circle = 2(pi)(r) Area of a square = length of side^2 Circumference of a square = length of side*4 Circle C in Inches Radius Area 1 0.16 0.08 2 0.32 0.32 3 0.48 0.72 4 0.64 1.27 5 0.80 1.99 6 0.95 2.86 7 1.11 3.90 8 1.27 5.09 9 1.43 6.45 10 1.59 7.96 11 1.75 9.63 12 1.91 11.46 13 2.07 13.45 14 2.23 15.60 15 2.39 17.90 16 2.55 20.37 17 2.71 23.00 18 2.86 25.78 19 3.02 28.73 20 3.18 31.83 21 3.34 35.09 22 3.50 38.52 23 3.66 42.10 24 3.82 45.84 25 3.98 49.74 26 4.14 53.79 27 4.30 58.01 28 4.46 62.39 29 4.62 66.92 30 4.77 71.62 31 4.93 76.47 32 5.09 81.49 33 5.25 86.66 34 5.41 91.99 35 5.57 97.48 36 5.73 103.13 37 5.89 108.94 38 6.05 114.91 39 6.21 121.04 40 6.37 127.32 41 6.53 133.77 42 6.68 140.37 43 6.84 147.14 44 7.00 154.06 45 7.16 161.14 46 7.32 168.39 47 7.48 175.79 48 7.64 183.35 49 7.80 191.07 50 7.96 198.94 51 8.12 206.98 52 8.28 215.18 53 8.44 223.53 54 8.59 232.05 55 8.75 240.72 56 8.91 249.55 57 9.07 258.55 58 9.23 267.70 59 9.39 277.01 60 9.55 286.48 Square C in Inches Side Length Area 1 0.25 0.06 2 0.50 0.25 3 0.75 0.56 4 1.00 1.00 5 1.25 1.56 6 1.50 2.25 7 1.75 3.06 8 2.00 4.00 9 2.25 5.06 10 2.50 6.25 11 2.75 7.56 12 3.00 9.00 13 3.25 10.56 14 3.50 12.25 15 3.75 14.06 16 4.00 16.00 17 4.25 18.06 18 4.50 20.25 19 4.75 22.56 20 5.00 25.00 21 5.25 27.56 22 5.50 30.25 23 5.75 33.06 24 6.00 36.00 25 6.25 39.06 26 6.50 42.25 27 6.75 45.56 28 7.00 49.00 29 7.25 52.56 30 7.50 56.25 31 7.75 60.06 32 8.00 64.00 33 8.25 68.06 34 8.50 72.25 35 8.75 76.56 36 9.00 81.00 37 9.25 85.56 38 9.50 90.25 39 9.75 95.06 40 10.00 100.00 41 10.25 105.06 42 10.50 110.25 43 10.75 115.56 44 11.00 121.00 45 11.25 126.56 46 11.50 132.25 47 11.75 138.06 48 12.00 144.00 49 12.25 150.06 50 12.50 156.25 51 12.75 162.56 52 13.00 169.00 53 13.25 175.56 54 13.50 182.25 55 13.75 189.06 56 14.00 196.00 57 14.25 203.06 58 14.50 210.25 59 14.75 217.56 60 15.00 225.00

OpenStudy (anonymous):

Maximum Combined Area = 59 Inches of wire for Circle = 277.07 inches of Area Minimum Combined Area = 26 Inches of wire for Circle = 126.04 inches of Area